The reaction A--> B is first order in A. Initially [A]= 0.800M and after 54 min., [A]=0.100M. At what time (in min.) is [A]=(2.0e-2)M?
Answer #1

[A0]= 0.800M

[A]=0.100M

t = 54 min

K = 2.303/t log[A0]/[A]

K = 2.303/54 log0.8/0.1

    = 2.303*0.9030/54   = 0.0385 min-1

K = 2.303/t log[A0]/[A]

[A] = 0.02M

0.0385   = 2.303/t log0.8/0.02

0.0385    = 2.303*1.6020/t

         t= 2.303*1.6020/0.0385   = 95.82 min >>>>> answer